Kropf Inc. has provided the following data concerning one of the products in its standard cost system. Variable manufacturing overhead is applied to products on the basis of direct labor-hours.

Inputs

Standard Quantity or Hours per Unit of Output

Standard Price or Rate

Direct materials

8.90

liters

$

8.50

per liter

Direct labor

0.50

hours

$

29.70

per hour

Variable manufacturing overhead

0.50

hours

$

7.40

per hour

The company has reported the following actual results for the product for September:

Actual output

11,100

units

Raw materials purchased

100,500

liters

Actual cost of raw materials purchased

$

875,500

Raw materials used in production

98,820

liters

Actual direct labor-hours

5,140

hours

Actual direct labor cost

$

160,302

Actual variable overhead cost

$

33,414

Required:

a. Compute the materials price variance for September.

b. Compute the materials quantity variance for September.

c. Compute the labor rate variance for September.

d. Compute the labor efficiency variance for September.

e. Compute the variable overhead rate variance for September.

f. Compute the variable overhead efficiency variance for September.
Answer #1

a) Material price variance = (8.5*100500-875500) = 21250 U

b) Material quantity variance = (11100*8.9-98820)*8.5 = 255 U

c) Labor rate variance = (29.70*5140-160302) = 7644 U

d) Labor efficiency variance = (5550-5140)*29.70 = 12177 F

e) Variable overhead rate variance = (7.4*5140-33414) = 4622 F

f) Variable overhead efficiency variance = (5550-5140)*7.4 = 3034 F